OpenStudy (anonymous):

Factor the expression: x^2-x-12

OpenStudy (anonymous):

I put (x+4)(x-3) and got it wrong.

OpenStudy (jadeishere):

That's because it's (x - 4)(x + 3) x^2 + 3x - 4x - 12 It's the negative sign that has to be in the right spot :) the 'x' has to be a negative and -3 + 4 = 1 -4 + 3 = -1